The city council wants to sponsor 150 bags for their organization’s scholars. Each bag


costs $375. If they were to buy from the factory, they will be given 45% off the total


purchase. However, if they buy from the store they will be given 50% off the price for


each bag. Where will they have a cheaper purchase?


1
Expert's answer
2022-03-27T07:44:37-0400

Number of bags -150

Each bag cost-$375

"total cost =375×150=\$56250"

For factory purchase

"Discount=45\%"

"\frac{45}{100}=\frac{9}{20}=\frac{x}{56250}"

"x=25312.5"

"56250-25312.5=\$30937.5"


For store purchase

"Discount=50\%"

"\frac{50}{100}=\frac{1}{2}=\frac{x}{56250}"

"x=28125"

"56250-28125=\$28125"


At store they will have a cheaper purchase of "\$28125"